# Vision-Language-Action Models: Concepts, Progress, Applications and Challenges

### v1

@@ 3.4.5 Precision and Automated Agriculture
As illustrated in Figure 15, VLA models are emerging as transformative tools in precision and automated agriculture, offering intelligent, adaptive solutions for labor-intensive tasks across diverse farming landscapes [57, 150]. Unlike traditional agricultural automation systems that depend on rigid, sensor-driven pipelines and require manual reprogramming for each task or environmental variation [169, 84], VLAs integrate multimodal perception, natural language understanding, and real-time action generation within a unified framework [131, 66]. This enables autonomous ground robots and drones to interpret complex field scenes, follow spoken or text-based farming instructions, and generate context-aware actions such as selective fruit picking or adaptive irrigation. The ability of VLAs to dynamically adjust to occlusions, terrain irregularities, or varying crop types—combined with training on synthetic, photorealistic datasets—allows them to generalize across geographies and seasons. By leveraging action tokenization [189], transformer-based policy generation [11, 67], and techniques like LoRA fine-tuning [72], these systems are redefining the scalability and intelligence of agricultural robotics for sustainable and precision-driven farming.
Figure 15: This diagram illustrates the application of VLA models in precision and automated agriculture. A ground robot uses vision encoders to detect ripe fruits and interprets instructions such as “pick only Grade A fruits” through language encoders. Action tokens then guide robotic manipulators for efficient, damage-free picking. Drones leverage VLA models to analyze aerial imagery and verbal commands for targeted irrigation. Synthetic training environments and LoRA-based adaptation enable models to generalize across crop types, environmental conditions, and geographies. This VLA-driven pipeline promotes sustainable agriculture by improving productivity, reducing manual labor, and enhancing decision-making through multimodal perception and control.
@@
In modern orchards and crop fields, VLAs can process visual inputs from RGB-D cameras, multispectral sensors, or drones to monitor plant growth, detect diseases, and identify nutrient deficiencies. Vision transformers (e.g., ConvNeXt, DINOv2) encode spatial and semantic information from visual scenes, while large language models (e.g., T5, LLaMA) parse natural language commands—such as “inspect the east plot for powdery mildew” or “harvest ripe apples near the irrigation trench.” Through token fusion, these modalities are aligned in a shared representation space, allowing robots to execute fine-grained, context-aware actions with precision.
@@
For instance, in fruit-picking tasks, as illustrated in Figure 15, a VLA-equipped ground robot can identify ripe produce using image-based ripeness cues, interpret user-specified criteria such as “pick only Grade A fruits,” and execute motion sequences via action tokens that control its end-effector. This approach ensures minimal crop damage, optimizes pick rates, and allows real-time adaptation to unexpected variables like occlusions or terrain shifts. In irrigation management, drones guided by VLA models can interpret field maps and verbal instructions to selectively water stressed zones, reducing water usage by up to 30%.
Moreover, VLA models support dynamic reconfiguration and lifelong learning. With access to synthetic training datasets generated from photorealistic simulations of crop environments (e.g., 3D orchard renderings), models can be trained to recognize pests, weeds, and crop maturity stages without extensive manual annotation. Techniques like LoRA adapters and diffusion-based policy tuning further enhance generalization to novel crops, seasons, and geographical regions.
@@
The integration of VLAs into agricultural workflows offers significant benefits: reduced dependence on skilled labor, higher yield through targeted intervention, and enhanced environmental sustainability through optimized input usage. As global food systems grapple with climate variability and resource constraints, VLA-enabled agriculture will play a pivotal role in advancing scalable, intelligent, and sustainable farming practices tailored to real-world complexity.
@@ 3.4.6 Interactive AR Navigation with Vision-Language-Action Models
Interactive Augmented Reality (AR) navigation represents a frontier where the VLA models can significantly enhance human-environment interaction by providing intelligent, context-aware guidance in real-time [26, 80, 197]. In this paradigm, VLAs process continuous streams of visual data from AR-enabled devices—such as smart glasses or smartphones—alongside natural language queries to generate dynamic navigational cues overlaid directly onto the user’s view of the physical world. Unlike traditional GPS-based systems that rely on rigid maps and limited user input [24, 159], VLA-based AR agents interpret complex visual scenes (e.g., intersections, indoor hallways, signage) and respond to free-form instructions such as “take me to the nearest pharmacy with a wheelchair ramp” or “show the quietest route to the conference room.”
Technically, these models integrate a vision encoder (e.g., ViT, DINOv2) that extracts scene representations from first-person RGB frames, a language encoder (e.g., T5 or LLaMA) that processes user prompts or voice commands, and an action decoder that predicts tokenized navigation cues such as directional overlays, waypoints, or voice instructions. A transformer-based architecture fuses these modalities to reason about both the spatial layout and semantic intent, allowing the AR agent to adaptively highlight paths, landmarks, and hazards directly within the user’s field of view [163, 129]. For example, as shown in Figure 16, in a crowded airport, the VLA agent could visually identify escalators, gates, or baggage claims while understanding a query like “how do I reach Gate 22 without stairs?”, adjusting the route in response to real-time occupancy and obstacles.
VLAs also support interaction loops that enable users to refine instructions (e.g., “avoid busy areas” or “take the scenic route”) and receive context-aware feedback, improving accessibility for the visually impaired or cognitively challenged. In logistics and indoor navigation, these systems can be integrated with IoT sensors and digital twins to guide warehouse workers, maintenance teams, or delivery robots through complex environments. Furthermore, personalized navigation can be achieved through continual fine-tuning, where VLA models learn user preferences and local spatial layouts over time.
@@
As AR hardware becomes more affordable and integrated into daily life, VLA-powered navigation systems will enable seamless spatial understanding, multimodal interaction, and autonomous guidance in public, industrial, and assistive contexts—redefining how humans perceive, explore, and interact with physical spaces.
@@ 4 Challenges and Limitations of Vision-Language-Action Models
VLA models face a spectrum of interrelated challenges that impede their translation from research prototypes to robust, real‐world systems. First, achieving real‐time, resource‐aware inference remains difficult: models like DeeR-VLA leverage dynamic early‐exit architectures to cut computation 5–6× on manipulation benchmarks while preserving accuracy, yet their gains diminish in complex scenarios [202]. Similarly, Uni-NaVid compresses egocentric video tokens for 5 Hz navigation but still struggles under highly ambiguous instructions and longer horizons [210]. Coupled with limited object generalization, even advanced hybrid vision‐language grounding schemes (e.g., ObjectVLA) generalize to only 64 % of novel objects, underscoring persistent gaps in open‐world robustness [223].
@@
Second, adapting VLA models with minimal supervision and ensuring stable policy updates under scarce, noisy data is nontrivial. ConRFT combines behavior cloning and Q-learning with human‐in‐the‐loop fine‐tuning to rapidly converge to 96.3% success over eight contact‐rich tasks, yet it relies heavily on expert interventions and reward shaping [31]. Hierarchical frameworks such as Hi Robot decouple high‐level reasoning from low‐level execution to improve instruction fidelity, but coordinating these modules and grounding ambiguous feedback remains challenging [155]. Likewise, TLA’s fusion of tactile streams with language commands achieves over 85 % success on unseen peg‐in-hole tasks, but dataset breadth and real‐time multi‐step decoding still limit broader generalization [70].
@@
Furthermore, ensuring safety, generalization, and end‐to‐end reliability in dynamic environments demands new modeling and evaluation standards. Occupancy‐Language‐Action models like OccLLaMA unify 3D scene understanding with action planning, yet they must scale to richer scene dynamics and semantic consistency across modalities [183]. RaceVLA pushes high‐speed drone navigation via quantized, iterative control loops, but its visual–physical generalization trails larger VLAs and dedicated reasoners [153]. Model‐merging strategies in ReVLA recover lost out‐of‐domain visual robustness—improving OOD grasp success by up to 77 %—but introduce extra computation and complexity [39]. Finally, SafeVLA formulates constraints via constrained Markov decision processes to cut unsafe behavior by over 80 %, yet defining comprehensive, non‐restrictive safety rules for diverse real‐world tasks remains an open problem [205]. Addressing these intersecting limitations is critical for VLA models to achieve reliable, autonomous operation against the full complexity of real‐world robotics.
@@
Building upon the critical limitations outlined above, it is imperative to map each challenge to targeted mitigation strategies and forecast their system-level impact. Table 4 distills this mapping into three columns—identifying core limitations, proposing concrete technical remedies drawn from recent advances, and articulating the anticipated benefits for real-world VLA deployment. For instance, tackling real-time inference constraints leverages parallel decoding and quantized transformer pipelines with hardware acceleration (e.g., TensorRT) to sustain control loop rates in drones and manipulators [100, 94, 60, 110]. Addressing multimodal action representation via hybrid diffusion–autoregressive policies enriches a model’s capacity to produce varied, context-sensitive motor commands for complex tasks [133, 121]. To guarantee safety in open worlds, dynamic risk assessment modules and adaptive planning layers can be integrated, ensuring robust emergency stop behaviors in unpredictable settings [143, 180, 87]. Similarly, dataset bias and grounding are countered through curated debiased corpora and advanced contrastive fine-tuning, bolstering fairness and semantic fidelity when generalizing to novel objects and scenes [145, 16, 136]. Together, these solution pathways—and others spanning simulation-to-real transfer, tactile integration, and energy-efficient architectures—frame a comprehensive roadmap for transitioning VLA research into reliable, scalable autonomy.
Table 4: Challenges, Potential Solutions, and Expected Impact of VLA Models
@@
Real-Time Inference Constraints Adopt parallel decoding, quantized transformers, and hardware acceleration [100, 94](e.g., TensorRT); minimize autoregressive overhead [60, 110] Supports real-time robotic control and broader deployment in time-sensitive domains[194, 150] (e.g., drones, manipulators)
@@
Multimodal Action Representation Hybrid tokenization using diffusion and autoregressive policies [133]; train on diverse demonstrations and multi-modal outputs [121] Improves handling of complex, dynamic manipulation tasks with multiple viable solutions [59]
@@
Safety Assurance in Open Worlds Integrate dynamic risk assessment modules [143, 180]; low-latency emergency stop circuits and adaptive planning layers [87] Ensures reliability and safety in unpredictable environments (homes, factories, healthcare settings)
@@
Dataset Bias and Grounding Curate diverse, debiased datasets [145]; apply improved grounding techniques such as CLIP fine-tuning with hard negatives [212, 16] Enhances model fairness, semantic fidelity [85], and generalizability to novel real-world inputs [175, 217, 136]
@@
Limited 3D Perception and Reasoning Integrate 3D sensors (e.g., depth, LiDAR), develop 3D-aware architectures, and leverage point cloud fusion with vision-language inputs Enables more accurate spatial reasoning, manipulation, and navigation in complex real-world environments [100]
@@
Cross-Embodiment Generalization Train with diverse robot types and morphologies, use embodiment-agnostic representations, and apply cross-domain adaptation techniques [201] Facilitates transfer of policies and knowledge across different robot platforms and configurations [209, 94]
@@
Annotation Complexity and Cost Employ weak supervision, active learning, and synthetic data generation to reduce reliance on extensive manual annotation [115] Lowers development costs and accelerates scaling to new tasks and domains [180, 215]
@@
Sim-to-Real Transfer Gap Use domain adaptation, sim-to-real fine-tuning, and real-world calibration strategies[162, 104] Improves reliability and consistency of VLA models when deployed outside simulation environments [4, 53]
@@
Integration of Physical Knowledge Incorporate physics-based priors, simulation environments, and dynamics modeling into training pipelines [42] Enhances the model’s ability to predict and plan actions that respect real-world physical constraints [82]
@@
Multi-Modal Integration (e.g., tactile, audio) Fuse additional sensory modalities (tactile, audio) with vision and language [88]; develop multi-modal transformers Expands task repertoire and robustness to ambiguous or visually occluded scenarios [61, 106, 71]
@@
Handling Long-Horizon, Multi-Stage Tasks Design hierarchical policies, memory-augmented networks, and trajectory planning modules [106] Improves performance on complex, sequential tasks requiring planning and memory [102, 215, 175, 136]
@@
System Integration Complexity Develop unified Transformer backbones [218]; incorporate temporal alignment layers and sim-to-real transfer learning strategies [127, 207] Enables seamless planning-control coordination and robust transfer to physical robots [147, 160]
@@
Energy and Compute Demands Apply model pruning, LoRA adapters, quantization-aware training, and deployment on low-power accelerators Facilitates scalable, efficient deployment of VLAs in embedded and mobile platforms [195, 213, 96, 190]
@@
Generalization to Unseen Tasks Use compositional generalization, few-shot meta-learning, and task-agnostic pretraining pipelines [135, 113] Reduces task-specific overfitting, enabling robust zero- and few-shot adaptation [75, 186, 215]
Robustness to Environmental Variability Use domain randomization, sensor fusion, and real-time recalibration of perception-action pipelines [121] Enhances performance in changing or cluttered environments with minimal degradation [214, 186]
@@
Ethical and Societal Implications Enforce privacy via on-device processing and anonymization [116, 154, 195, 29]; audit model fairness; build regulatory frameworks for trust Promotes equitable and trustworthy VLA adoption across social, medical, and labor domains [124, 140, 166, 134]
@@
The remainder of this section is organized into five focused subsections, each examining a distinct cluster of VLA challenges identified in the literature. First, we analyze real-time inference constraints and the emerging methods that address them. Next, we delve into multimodal action representation alongside safety assurance in open-world settings. We then discuss dataset bias, grounding strategies, and generalization to unseen tasks, followed by an exploration of system integration complexity and computational demands. Finally, we consider robustness and the ethical implications of deploying VLAs in real-world applications.

### v2

@@ 3.4.5 Precision and Automated Agriculture
As illustrated in Figure 15, VLA models are emerging as transformative tools in precision and automated agriculture, offering intelligent, adaptive solutions for labor-intensive tasks across diverse farming landscapes [70, 191]. Unlike traditional agricultural automation systems that depend on rigid, sensor-driven pipelines requiring manual reprogramming for each task or environmental variation [220, 108], VLAs integrate multi-modal perception, natural language understanding, and real-time action generation within a unified framework [167, 84]. This unified multimodal integration enables autonomous ground robots and drones to interpret complex field scenes, follow spoken or text-based farming instructions, and generate context-aware actions such as selective fruit picking or adaptive irrigation. The ability of VLAs to dynamically adjust to occlusions, terrain irregularities, lighting variability, or varying crop types combined with training on synthetic, photorealistic datasets allows them to generalize across crop types, geographies and seasons. By leveraging action tokenization [245], transformer-based policy generation [12, 85], and techniques like LoRA fine-tuning [93], these systems are redefining the scalability and intelligence of agricultural robotics for sustainable and precision-driven farming.
Figure 15: Conceptual illustration of VLA models in precision and automated agriculture. A ground robot combines vision encoders and language instructions (e.g., “pick only Grade A fruits”) to generate action tokens for damage-free harvesting, while aerial robots use VLA reasoning for language-guided irrigation. Synthetic training, LoRA-based adaptation, and lifelong feedback enable generalization across crops, environments, and geographies, supporting sustainable, data-driven agricultural automation.
@@
In modern fruit orchards and other crop fields, VLAs can process visual inputs from RGB-D cameras, multispectral sensors, or drones to monitor plant growth, detect diseases, and identify nutrient deficiencies. Vision transformers (e.g., ConvNeXt, DINOv2) encode spatial and semantic information from visual scenes, while LLMs (e.g., T5, LLaMA) parse natural language commands such as “inspect the east plot for powdery mildew” or “harvest ripe apples near the irrigation trench.” Through token fusion, these modalities are aligned in a shared representation space, allowing robots to execute fine-grained, context-aware actions with precision.
@@
For instance, in fruit-picking tasks, as illustrated in Figure 15, a VLA-equipped ground robot can identify ripe produce using image-based ripeness cues, interpret user-specified criteria such as “pick only Grade A fruits,” and execute motion sequences via action tokens that control its end-effector. This approach ensures minimal crop damage, optimizes pick rates, and allows real-time adaptation to unexpected variables like occlusions or terrain shifts. In irrigation management, drones guided by VLA models can interpret field maps and verbal instructions to selectively water stressed zones, reducing water usage %.
Beyond immediate task execution, VLA models are expected to support dynamic reconfiguration and lifelong learning through closed-loop feedback mechanisms. In particular, execution outcomes, sensory observations, and task success signals collected during deployment can be logged and periodically incorporated into offline or incremental updates of the VLA policy. When combined with synthetic training data generated from photorealistic simulations of crop environments (e.g., 3D orchard renderings), such feedback-driven adaptation may enable models to progressively improve robustness to new crop varieties, pest conditions, and seasonal variations without extensive manual annotation. Parameter-efficient techniques such as LoRA adapters and diffusion-based policy refinement are anticipated to play a key role in facilitating this continual adaptation while limiting computational overhead.
@@
Overall, the integration of VLA models into agricultural workflows is expected to offer several long-term benefits, including reduced reliance on skilled manual labor, improved yields through targeted intervention, and enhanced environmental sustainability via optimized input usage. As global food systems increasingly face climate variability and resource constraints, VLA-enabled agricultural technologies are anticipated to contribute to scalable, intelligent, and context-aware farming practices that better accommodate real-world complexity.
@@ 3.4.6 Interactive AR Navigation with Vision-Language-Action Models
Interactive Augmented Reality (AR) navigation represents a frontier where VLA models can significantly enhance human-environment interaction by providing intelligent, context-aware guidance in real-time [31, 104, 255]. In this paradigm, VLAs process continuous streams of visual data from AR-enabled devices such as smart glasses or smartphones alongside natural language queries to generate dynamic navigational cues overlaid directly onto the user’s view of the physical world. Unlike traditional GPS-based systems that rely on rigid maps and limited user input [29, 207], VLA-based AR agents interpret complex visual scenes (e.g., intersections, indoor hallways, signage) and respond to free-form instructions such as “take me to the nearest pharmacy with a wheelchair ramp” or “show the quietest route to the conference room.”
Technically, these models integrate a vision encoder (e.g., ViT, DINOv2) that extracts scene representations from RGB camera frames, a language encoder (e.g., T5 or LLaMA) that processes user prompts or voice commands, and an action decoder that predicts tokenized navigation cues such as directional overlays, waypoints, or voice instructions. A transformer-based architecture fuses these modalities to reason about both the spatial layout and semantic intent, allowing the AR agent to adaptively highlight paths, landmarks, and hazards directly within the user’s field of view [211, 165]. For example, as shown in Figure 16, in a crowded airport, the VLA agent could visually identify escalators, gates, or baggage claims while understanding a query like “how do I reach Gate 22 without stairs?”, adjusting the route in response to real-time occupancy and obstacles.
VLAs are expected to support interactive instruction loops in which users can first issue high-level commands (e.g., “navigate to the pharmacy”) and subsequently refine them with additional constraints such as “avoid busy areas” or “take the scenic route.” Through context-aware feedback and iterative clarification, such interaction paradigms can improve accessibility and usability for visually impaired or cognitively challenged individuals. In logistics and indoor navigation, these systems can be integrated with IoT sensors and digital twins to guide warehouse workers, maintenance teams, or delivery robots through complex environments. Furthermore, personalized navigation can be achieved through continual fine-tuning, where VLA models learn user preferences and local spatial layouts over time.
@@
As AR hardware becomes more affordable and integrated into daily life, VLA-powered navigation systems will enable seamless spatial understanding, multi-modal interaction, and autonomous guidance in public, industrial, and assistive contexts redefining how humans perceive, explore, and interact with physical spaces.
@@ 4 Challenges and Limitations of Vision-Language-Action Models
VLA models face a spectrum of interrelated challenges that impede their translation from research prototypes to robust, real‐world systems. First, achieving real‐time, resource‐aware inference remains difficult: models like DeeR-VLA leverage dynamic early‐exit architectures to cut computation 5–6× on manipulation benchmarks while preserving accuracy, yet their gains diminish in complex scenarios [269]. Similarly, Uni-NaVid compresses egocentric video tokens for 5 Hz navigation but still struggles under highly ambiguous instructions and longer horizons [280]. Moreover, these efficiency-driven designs often expose a trade-off between computational speed and representational coverage. When operating under aggressive compression or early-exit constraints, even advanced hybrid vision–language grounding methods exhibit limited object generalization; for example, ObjectVLA generalizes to only 64 % of novel objects, highlighting how real-time optimization can exacerbate gaps in open-world robustness [297].
@@
Second, adapting VLA models with minimal supervision and ensuring stable policy updates under scarce, noisy data is nontrivial. ConRFT combines behavior cloning and Q-learning with human‐in‐the‐loop fine‐tuning to rapidly converge to 96.3% success over eight contact‐rich tasks, yet it relies heavily on expert interventions and reward shaping [36]. Hierarchical frameworks such as Hi Robot decouple high‐level reasoning from low‐level execution to improve instruction fidelity, but coordinating these modules and grounding ambiguous feedback remains challenging [199]. Likewise, Tactile-language-action model model’s fusion of tactile streams with language commands achieves over 85 % success on unseen peg‐in-hole tasks, but dataset breadth and real‐time multi‐step decoding still limit broader generalization [88].
@@
Furthermore, ensuring safety, generalization, and end‐to‐end reliability in dynamic environments demands new modeling and evaluation standards. Occupancy‐Language‐Action models like OccLLaMA unify 3D scene understanding with action planning, yet they need to scale to richer scene dynamics and semantic consistency across modalities [239]. RaceVLA advances high-speed drone navigation through quantized, iterative control loops; however, its limited visual–physical generalization compared to larger VLAs and dedicated reasoning models raises safety concerns in unseen or rapidly changing environments [195]. Model‐merging strategies in ReVLA recover lost out‐of‐domain visual robustness improving oriented object detection (OOD) grasp success by up to 77 % but introduce extra computation and complexity [48]. Finally, SafeVLA formulates constraints via constrained Markov decision processes to cut unsafe behavior by over 80 %, yet defining comprehensive, non‐restrictive safety rules for diverse real‐world tasks remains an open problem [274]. Addressing these intersecting limitations is critical for VLA models to achieve reliable, autonomous operation under the full complexity of real‐world robotics.
@@
Building upon the critical limitations outlined above, it is imperative to map each challenge to targeted mitigation strategies and assess their system-level impact. Table 4 presents this map identifying core limitations, potential technical remedies drawn from recent advances, and articulating the anticipated benefits for real-world VLA deployment. For instance, addressing real-time inference constraints leverages parallel decoding and quantized transformer pipelines with hardware acceleration (e.g., TensorRT) to sustain control loop speed in drones and manipulators [129, 122, 75, 142]. Addressing multi-modal action representation via hybrid diffusion–autoregressive policies enriches a model’s capacity to produce varied, context-sensitive motor commands for complex tasks [171, 156]. To guarantee safety in open worlds, dynamic risk assessment modules and adaptive planning layers can be integrated, ensuring robust emergency stop behaviors in unpredictable settings [183, 233, 113]. Similarly, dataset bias and grounding can be minimized through curated debiased corpora and advanced contrastive fine-tuning, strengthening fairness and semantic fidelity when generalizing to novel objects and scenes [185, 17, 175]. Together, these strategies and other approaches spanning simulation-to-real transfer, tactile integration, and energy-efficient architectures frame a comprehensive roadmap for transitioning VLA research into reliable, scalable autonomy.
Table 4: Challenges, potential solutions, and expected impact of VLA models.
@@
Real-time inference constraints Parallel decoding, quantized transformers, and hardware acceleration (e.g., TensorRT) [129, 122]; reduce autoregressive overhead [75, 142]. Enables real-time control and deployment in latency-critical domains [251, 191] (e.g., UAVs, manipulators).
@@
Multi-modal action representation Hybrid tokenization combining diffusion and autoregressive policies [171]; train on diverse demonstrations and multi-modal outputs [156]. Improves performance on complex, dynamic manipulation with multiple valid solution modes [74].
@@
Safety assurance in open worlds Dynamic risk assessment modules [183, 233]; low-latency emergency-stop and adaptive planning layers [113]. Improves reliability and safety in unpredictable settings (homes, factories, healthcare); enhances user acceptability.
@@
Dataset bias and grounding Curate diverse/debiased datasets [185]; stronger grounding (e.g., CLIP fine-tuning with hard negatives) [282, 17]. Improves fairness and semantic fidelity [109], and enhances generalization to novel real-world inputs [227, 288, 175].
@@
Limited 3D perception and reasoning Integrate depth/LiDAR; develop 3D-aware architectures; fuse point clouds with vision–language features. Enables stronger spatial reasoning for manipulation and navigation in complex environments [129].
@@
Cross-embodiment generalization Train across diverse morphologies; learn embodiment-agnostic action abstractions; apply cross-domain adaptation [266]. Facilitates policy transfer across robot platforms and configurations [279, 122].
@@
Annotation complexity and cost Weak supervision, active learning, and synthetic data generation to reduce manual labeling [148]. Lowers development cost and accelerates scaling to new tasks/domains [233, 286].
@@
Sim-to-real transfer gap Domain adaptation, sim-to-real fine-tuning, and real-world calibration [210, 135]. Improves reliability and consistency when deploying beyond simulation [4, 66].
@@
Integration of physical knowledge Physics priors, simulation environments, and dynamics modeling in training pipelines [53]. Improves prediction and planning under real physical constraints [106].
@@
Multi-modal integration (tactile, audio) Fuse tactile/audio with vision and language [114]; extend multimodal transformer fusion. Improves robustness under occlusion/ambiguity and expands the task repertoire [76, 136, 91].
@@
Long-horizon multi-stage tasks Hierarchical policies, memory-augmented networks, and trajectory planning modules [136]. Improves sequential planning, memory, and compositional execution [131, 286, 227, 175].
@@
System integration complexity Unified transformer backbones [289]; temporal alignment and sim-to-real transfer strategies [163, 276]. Enables tighter planning–control coordination and more robust transfer to physical robots [187, 208].
@@
Energy and compute demands Pruning, LoRA, quantization-aware training, and low-power accelerators. Enables efficient embedded/mobile deployment [252, 283, 125, 246].
@@
Generalization to unseen tasks Compositional generalization, few-shot meta-learning, and task-agnostic pretraining [173, 147]. Reduces overfitting and strengthens zero-/few-shot adaptation [97, 242, 286].
Robustness to environmental variability Domain randomization, sensor fusion, broader training datasets, and online recalibration [156]. Improves stability under changing lighting, clutter, and scene dynamics [285, 242].
@@
Ethical and societal implications Privacy via on-device processing/anonymization [149, 198, 252, 34]; fairness audits; regulatory and trust frameworks. Promotes equitable and trustworthy adoption across social, medical, and labor domains [160, 180, 217, 172].
@@
The remainder of this section is organized into five focused subsections, each examining a distinct cluster of VLA challenges identified in the literature. First, we analyze real-time inference constraints and the emerging methods to address them. Next, we explore multi-modal action representation alongside safety assurance in open-world settings. We then discuss dataset bias, grounding strategies, and generalization to unseen tasks, followed by an exploration of system integration complexity and computational demands. Finally, we consider robustness and the ethical implications of deploying VLAs in real-world applications.